three times as many units of plywood is sold as lumber, the profit on the lumber is $25 per unit and the profit on the plywood is $40 per unit could you show me how to find how many units of lumber and plywood would be produced and sold to make a profit of $14,790
Answer by checkley77(12844)

3X*40+X*25=14,790
120X+25X=14,790
145X=14,790
X=14,790/145
X=102 PIECES OF LUMBER WAS SOLD.
3*102=306 PIECES OF PLYWOOD WERE SOLD.
3*102*40+102*25=14,790
12,240+2,550=14,790
14,790=14,790